OpenStudy (anonymous):

Simplify. The variable represents a nonnegative real number. √147b^3/b

OpenStudy (anonymous):

$$ 7 \sqrt{3} b^2 $$

OpenStudy (anonymous):

Is it \[7\sqrt{3b ^{2}} or 7\sqrt{3}b ^{2}\]

OpenStudy (anonymous):

$$ \large 7 \sqrt{3} \times b^2$$